# AlphaForge: A Framework to Mine and Dynamically Combine Formulaic Alpha Factors

### v1

@@ 4.1.3. Additional Experiments
4.2. Evaluation Metrics
IC: The time-averaged cross-sectional Pearson correlation between factor values and actual returns, as expressed in Equation 1.
ICIR: A metric that considers the stability of an alpha factor, as shown in Equation 2 .
RankIC: Ranking IC, as shown in Equation 3.
Rank ICIR: Ranking ICIR, as shown in Equation 4.
4.3. Main Results
Regarding Q1: Table 1 presents the comparative analysis. As shown in Table 1, our method demonstrates superior performance across various metrics, including stock selection ability indicators such as IC and RankIC, as well as stability indicators like ICIR and RankICIR. This indicates that the AlphaForge framework has achieved notable advancements in both stock selection ability and stability compared to the baseline methods. While the traditional GP approach exhibited limited performance, the reinforcement learning-based group Alpha factor generation method RL, displayed relatively stronger performance. On top of that, our method surpasses these comparison baseline models in terms of all metrics. This outcome proves the effectiveness of our framework.
Table 1. Results of the Metrics in csi300 and csi500. Standard deviations are in the brackets.
CSI300
Method IC Rank IC ICIR RankICIR
GP 1.29%(0.44%) 2.72%(0.58%) 0.072(0.026) 0.140(0.033)
RL 2.09%(0.26%) 2.72%(0.42%) 0.141(0.021) 0.167(0.027)
Ours 4.40%(0.56%) 5.89%(0.69%) 0.368(0.041) 0.454(0.060)
CSI500
Method IC Rank IC ICIR RankICIR
GP 0.37%(0.76%) 2.34%(1.07%) 0.024(0.046) 0.124(0.058)
RL 1.91%(0.49%) 4.03%(0.62%) 0.152(0.042) 0.286(0.056)
Ours 2.84%(0.58%) 5.57%(0.58%) 0.212(0.043) 0.386(0.042)
4.4. Effect of the Pool Size
Figure 3. The IC in CSI300 across Different Pool Size
@@ 4.5. Ablation Study
Table 2. Results of Ablation Study, standard deviations in the brackets.
CSI300
Method IC Rank IC ICIR RankICIR
RL 2.09%(0.26%) 2.72%(0.42%) 0.141(0.021) 0.167(0.027)
Static 2.43%(0.57%) 3.67%(0.46%) 0.144(0.036) 0.204(0.029)
Dynamic 4.40%(0.56%) 5.89%(0.69%) 0.368(0.041) 0.454(0.060)
CSI500
Method IC Rank IC ICIR RankICIR
RL 1.91%(0.49%) 4.03%(0.62%) 0.152(0.042) 0.286(0.056)
Static 2.05%(0.29%) 4.48%(0.46%) 0.180(0.040) 0.336(0.056)
Dynamic 2.84%(0.58%) 5.57%(0.58%) 0.212(0.043) 0.386(0.042)
4.6. Case Study
A case study of a composite model was extracted to illustrate our framework’s capability in dynamic factor timing. Our generative model produced a factor zoo with 100 alpha factors, while the factor pool limit of the composite model was set at 10. Tables 3 and 4 present the composition of the ”Mega-Alpha” factor and the corresponding weights on two distinct trading days, respectively. It is observed that, among the 10 alpha factors selected on the first trading day, only 5 are used on the second trading day. Notably, Factor 3 had a weight of -0.00014 on the first trading day, whereas its weight shifted to 0.00168 on the second trading day. This indicates that the same factor contributed differently to the ”Mega-Alpha” generated by our model on different dates, highlighting the effectiveness of our framework in leveraging diverse Alpha factors to generate signals on various dates. This underscores a process of timing selection for alpha factors within our framework.
Table 3. Factors used on Day 1
index exprs weight
@@ 4.6. Case Study
Table 4. Factors used on Day 2
index exprs weight
@@
4.7. Simulated Trading
@@
To assess the practical efficacy of our model, we conducted simulated trading based on the prediction results. The simulation trading period spanned from January 1, 2018, to December 31, 2022, utilizing the CSI300 stock pool. We employed the Qlib(Yang et al., 2020) framework, where the trading strategy entailed holding the top 50 stocks with the highest Mega-Alpha scores on a daily equal-weighted basis. Additionally, a daily limit of changing a maximum of 5 stocks was imposed to avoid excessive trading costs.
Figure 4. The simulated Trading Result
@@
By tracking the net value of funds in a simulated trading account. Figure 4 illustrates the cumulative returns of different algorithms. The results show that our framework is able to achieve the best net account value in a simulation trade lasting for five years. It performs the strongest among all the comparative models.
5. Related Work
@@
Formulaic Alpha Factors: The exploration of Alpha factor formulation encompasses a vast search space. Genetic programming algorithm has historically been employed to generate factors through tree-based genetic mutations. Early advancements, notably within the GPLearn package,(Lin et al., 2019a) introduced time series operators, establishing the first genetic programming method for mining alpha factors . (Lin et al., 2019b) employed mutual information as the objective for factor mining to discover Alpha factors based on nonlinear relationships. (Zhang et al., 2020) utilized the IC between Alphas to filter overly similar Alphas, enhancing the diversity of mined Alpha factors. AlphaEvolv (Cui et al., 2021) aimed at improving existing Alpha factors. Presently, the state-of-the-art method , as outlined in (Yu et al., 2023), is based on reinforcement learning (RL) for synergistic Alpha factor mining, offering a novel approach beyond genetic programming. This method leverages the space exploration capability of reinforcement learning to mine a set of collaborative Alphas, considering the synergies between highly correlated factors. However, current methods often fail to account for time-varying factor effects, typically adopting fixed factor combination weights and immutable objective functions during mining.
@@
Machine Learning-based Alpha Factors: Methods for predicting stock returns using deep learning are also thriving. Early approaches did not consider interactions between stocks, merely using the historical time series data of each stock to predict stock prices, including Multilayer Perceptron (MLP), Transformer (Vaswani et al., 2017), LSTM (Hochreiter and Schmidhuber, 1997), as well as tree-based methods like LightGBM (Ke et al., 2017) and XGBoost (Chen and Guestrin, 2016). Subsequent developments led to models specifically designed for this task, such as HIST (Xu et al., 2021a) which enhances traditional time series models by incorporating industry and concept graph data to capture the correlation information between different stocks. FactorVAE(Duan et al., 2022) combines the dynamic factor model (DFM) and variational autoencoder (VAE) to estimate the variance of the latent space distribution while predicting stock returns. These machine learning-based methods, when compared to the Formulaic Alpha approach, often lack interpretability.
@@
Discussion: Our framework introduces a novel approach for mining formulaic Alpha factors, with a primary focus on the stock investment domain. This framework could be expanded to many other domains including but not limited to traffic flow prediction(Ji et al., 2023), sales forecasting(Zhao et al., 2022; Zhang and Lv, 2023), and customer churn prediction(Khattak et al., 2023). Despite the optimization objective of our mining model changing with the increase of the factor zoo, the training objective of our framework’s mining algorithm’s generator is the IC of individual alpha factors. Exploring the possibility of incorporating the combined IC of the entire batch of factors may enhance the efficiency of factor mining.
6. Conclusion
@@
This paper introduces a new framework, named AlphaForge, for mining and dynamically combining formulaic alpha factors, thereby providing new tools for investors engaged in quantitative trading. Our AlphaForge framework is able to leverage the powerful space exploration capabilities of deep learning models to explore the search space for formulaic Alphas. The design of the surrogate model to predict Fitness Score ensures that our model can efficiently generate Alpha factors using gradient methods, even in scenarios characterized by a sparse search space. We also introduce a composite model capable of dynamically combining factor weights at each time slice, allowing the Mega-Alpha generated by the model to timely adjust its component factor and their weights to adapt to market fluctuations. Through extensive experiments, we have demonstrated that our framework can achieve better performance compared to previous methods of formulaic alpha models. In more realistic trading tests, our model consistently delivers higher profits.
Acknowledgements.
This work was supported by Beijing Natural Science Foundation under Grant IS23128 and the Ministry of Science and Technology of China under Grant 2020AAA0108404.
Appendix A Operators
We extended some operators based on (Yu et al., 2023), and Table 5 lists all the operators we provide. Here, C denotes operators for Cross-Section, which consider only the data of the current day in their calculations. T denotes time-series operators, which take into account data from a past period in their calculations. Unary and binary indicate the number of parameters required by the operator.
Table 5. Operators within Our Framework. C: cross-sectional, T: time-series
Operator Category Description
Abs(x) unary-C The absolute value $|x|$.
Log(x) unary-C Natural logarithmic function $\log(x)$.
[…]
Inv(x) unary-C The multiplicative inverse of the operand, represented as $1/x$.
S_log1p(x) unary-C Applies the signed log1p function, which is $\text{sign}(x)\cdot\log(1+|x|)$, to the operand for numerical stability.
CSRank(x) unary-C Computes the cross-sectional rank of elements within a tensor, normalized by the number of non-NaN elements and assigns NaN where the original tensor has NaN.
Greater(x, y)
Less(x, y) binary-C The larger/smaller one of the two values.
Power(x, y) binary-C Calculates the signed power of $x$ to the power of $y$, represented as $\text{sign}(x)\cdot|x|^{y}$.
Ref(x, t) unary-T The expression $x$ evaluated at $t$ days before the current day.
Mean(x, t)
Med(x, t)
Sum(x, t) unary-T The mean/median/sum value of the expression $x$ evaluated on the recent $t$ days.
Std(x, t)
Var(x, t) unary-T The standard deviation/variance of the expression $x$ evaluated on recent $t$ days.
Max(x, t)
Min(x, t) unary-T The maximum/minimum value of the expression $x$ evaluated on the recent $t$ days.
Mad(x, t) unary-T The mean absolute deviation $\mathrm{E}[|x-\mathrm{E}[x]|]$ of the expression $x$ evaluated on the recent $t$ days.
Delta(x, t) unary-T The relative difference of $x$ compared to $t$ days ago, $x-\text{Ref}(x,t)$.
WMA(x, t)
EMA(x, t) unary-T Weighted moving average and exponential moving average of the expression $x$ evaluated on the recent $t$ days.
Pctchange(x, t) unary-T Computes the percentage change in a time series, calculated as value at time $\frac{\text{value at time }t+\Delta t}{\text{value at time }t}-1$.
Skew(x, t) unary-T Calculates the skewness of a time series, defined as the third central moment divided by the cubed square root of the second central moment, $\frac{m3}{m2^{1.5}}$ in recent $t$ days.
Kurt(x, t) unary-T Calculates the kurtosis of a time series, computed as $\frac{m4}{\text{var}^{2}}-3$, measuring the ”tailedness” of the distribution in recent $t$ days.
IR(x, t) unary-T Calculates the Information Ratio of a time series, defined as the mean divided by the standard deviation of the series in recent $t$ days.
Min_Max_diff(x, t) unary-T Computes the difference between the maximum and minimum values of a time series over a specified window, highlighting the range within the dataset in recent $t$ days.
Cov(x, y, t) binary-T The covariance between two time series $x$ and $y$ in the recent $t$ days.
Corr(x, y, t) binary-T The Pearson’s correlation coefficient between two time series $x$ and $y$ in recent $t$ days.
Appendix B Hyper-parameters
The generative model employs DCGAN, consisting of three ConvTranspose2d layers with kernel sizes of [6,1], [5,1], and [6,1], along with three layers of 2d-CNN with a kernel size of [8,1]. The predictor utilizes 2D-CNN with kernel sizes of [1,3] and [1,4]. The batch size is set to 256, with a maximum formula length of 20, a latent space dimension of 100, and a dropout ratio of 0.2. The learning rate is 0.001. The latent space penalty coefficient is set to 10, and the similarity penalty coefficient is set to 1. $\operatorname{CORR}^{\prime}$ is set to be 0.5. The criterion for a factor to be included in the Factor Zoo is $IC>3\%$ and $ICIR>0.1$.

### v3

@@ Introduction
Additionally, due to space constraints, the user survey conducted among industry professionals and other supplementary experiments can be found in the supplementary materials.
Main Results
Regarding Q1: As shown in Table 1, our method demonstrates superior performance across various metrics, including stock selection ability indicators such as IC and RankIC. We conduct comparative experiments with non-formulative MLP,LlghtGBM, XGBoost, as well as with formulative methods GP, RL and DSO. Table 1 shows that our method outperforms all the methods in CSI300 and CSI500 datasets. This indicates that the AlphaForge framework has achieved notable advancements in stock selection ability compared to the baseline methods.
Table 1: Comparison of Methods on CSI 300 and CSI 500
CSI 300 CSI 500
IC(%) RankIC(%) IC(%) RankIC(%)
XGB 0.41 1.63 0.33 2.87
MLP 1.22(0.16) 1.75(0.28) 1.94(0.11) 3.31(0.23)
LGBM 0.84 1.85 1.75 3.81
GP 1.29(0.44) 2.72(0.58) 0.37(0.76) 2.34(1.07)
DSO 2.55(0.69) 3.88(1.12) 1.38(0.57) 4.56(0.61)
RL 2.09(0.26) 2.72(0.42) 1.91(0.49) 4.03(0.62)
Static 2.43 (0.57) 3.67(0.46) 2.05(0.29) 4.48(0.46)
Ours 4.40(0.56) 5.89(0.69) 2.84(0.58) 5.57(0.58)
Effect of the Pool Size
Figure 2: The IC in CSI300 across Different Pool Size
@@
Case Study
Table 2: Factors used on Day 1
# exprs weight
@@
Table 3: Factors used on Day 2
# exprs weight
@@
A case study of a composite model was extracted to illustrate our framework’s capability in dynamic factor timing. Our generative model produced a factor zoo with 100 alpha factors, while the factor pool limit of the composite model was set at 10. Tables 2 and 3 present the composition of the ”Mega-Alpha” factor and the corresponding weights on two distinct trading days, respectively. It is observed that, among the 10 alpha factors selected on the first trading day, only 5 are used on the second trading day. Notably, Factor 3 had a weight of -0.00014 on the first trading day, whereas its weight shifted to 0.00168 on the second trading day. This indicates that the same factor contributed differently to the ”Mega-Alpha” generated by our model on different dates, highlighting the effectiveness of our framework in leveraging diverse Alpha factors to generate signals on various dates. This underscores a process of timing selection for alpha factors within our framework.
Interpret-ability of Alpha Factors
Taking factor 1 in table 2 as an example, this factor can be interpreted as whether the trend of the high price and the volume are the same in the past 20 days. The negative weight of this factor reflects an underlying investment logic: When prices are rising but attract little attention, it may be worth considering a purchase. When prices are dropping and the crowd is panic selling, it may present an opportunity to buy.
Another example is the factor ‘-1*ts_mean(volume,20)’, which represents the opposite of average trading volume over the past 20 days. This factor has a strong correlation with the stock’s market capitalization. If the weight of this factor is too large in the model, it can cause the portfolio to lean towards small-cap stocks, leading to dangerous risk exposure. Usually, a competent investment manager should seek to reduce the weight of this factor or use other methods to avoid excessive exposure to small-cap stocks.
Investment managers could analyze the inherent meaning and return sources of each factor in the model and make adjustments based on their understanding of the market. Interpretability helps the investment manager to analyze the sources of returns and make adjustments and attributions in real-world investments based on their market insights.
Simulated Trading and Real Money Investment
@@
To assess the practical efficacy of our model, we conducted simulated trading based on the prediction results. The simulation trading period spanned from January 1, 2018, to December 31, 2022, utilizing the CSI300 stock pool. We employed the Qlib(Yang et al. 2020) framework, where the trading strategy entailed holding the top 50 stocks with the highest Mega-Alpha scores on a daily equal-weighted basis. Additionally, a daily limit of changing a maximum of 5 stocks was imposed to avoid excessive trading costs.
Figure 3: The Real(top) and simulated(bottom) Trading Result
@@
The top half shows our actual trading results. A real account was used with an investment of 3 million RMB in CSI500. Untill now, after approximately 9 months of investment, it has generated a 21.68% higher excess return compared to CSI500. By tracking the net value of funds in a simulated trading account. The bottom half of Figure 3 illustrates the cumulative returns of different algorithms. The results show that our framework is able to achieve the best net account value in a simulation trade lasting for five years. It performs the strongest among all the comparative models.
Related Work
@@
Formulaic Alpha Factors: The exploration of Alpha factor formulation encompasses a vast search space. Genetic programming algorithm has historically been employed to generate factors through tree-based genetic mutations. Early advancements, notably within the GPLearn package,(Lin et al. 2019a) introduced time series operators, establishing the first genetic programming method for mining alpha factors . (Lin et al. 2019b) employed mutual information as the objective for factor mining to discover factors based on nonlinear relationships. (Zhang et al. 2020a) utilized the IC between Alphas to filter overly similar Alphas, enhancing the diversity of mined Alpha factors. AlphaEvolv (Cui et al. 2021a) aimed at improving existing Alpha factors. Presently, (Yu et al. 2023) is based on reinforcement learning (RL) for synergistic Alpha factor mining, offering a novel approach beyond genetic programming. This method leverages the space exploration capability of reinforcement learning to mine a set of collaborative Alphas. However, current methods often fail to account for time-varying factor effects, typically adopting fixed factor combination weights and immutable objective functions during mining.
@@
Machine Learning-based Alpha Factors: Methods for predicting stock returns using deep learning are also thriving. Early approaches did not consider interactions between stocks, merely using the historical time series data of each stock to predict stock prices, including Multilayer Perceptron (MLP), Transformer (Vaswani et al. 2017a), LSTM (Hochreiter and Schmidhuber 1997), as well as tree-based methods like LightGBM (Ke et al. 2017) and XGBoost (Chen and Guestrin 2016). Subsequent developments led to models specifically designed for this task, such as HIST (Xu et al. 2021a) which enhances traditional time series models by incorporating industry and concept graph data to capture the correlation information between different stocks. FactorVAE(Duan et al. 2022) combines the dynamic factor model (DFM) and variational autoencoder (VAE) to estimate the variance of the latent space distribution while predicting stock returns. These machine learning-based methods, when compared to the Formulaic Alpha approach, often lack interpret-ability.
Symbolic regression addresses the problem of identifying relationships between different variables, typically aiming to derive a single interpretable mathematical expression to solve scientific problems. As one of the few interpretable machine learning methods, symbolic regression finds extensive applications in fields such as mathematics, physical equations, and materials science. Early approaches to symbolic regression focused on improving genetic programming (GP). For instance, (La Cava, Spector, and Danai 2016) used lexicase selection for regression, (Schmidt and Lipson 2010) introduced the age-layered population structure, and (Moraglio, Krawiec, and Johnson 2012) employed semantic variation operators to generate offspring. With the advancement of neural network technology, (Champion et al. 2019) proposed a new method by combining autoencoder networks with symbolic regression. (Biggio et al. 2021) emphasized the role of large-scale pre-training based on the transformer model, and (Landajuela et al. 2022a) introduced a unified framework of neural networks and symbolic regression.
@@
Discussion: Our framework introduces a novel approach for mining formulaic Alpha factors, with a primary focus on the stock investment domain. This framework could be expanded to many other domains including but not limited to traffic flow prediction(Ji et al. 2023), sales forecasting(Zhao, Zuo, and Lin 2022; Zhang and Lv 2023), and customer churn prediction(Khattak et al. 2023). Despite the optimization objective of our mining model changing with the increase of the factor zoo, the training objective of our framework’s mining algorithm’s generator is the IC of individual alpha factors. Exploring the possibility of incorporating the combined IC of the entire batch of factors may enhance the efficiency of factor mining.
Conclusion
@@
This paper introduces a new framework, named AlphaForge, for mining and dynamically combining formulaic alpha factors, thereby providing new tools for investors engaged in quantitative trading. Our AlphaForge framework is able to leverage the powerful space exploration capabilities of deep learning models to explore the search space for formulaic Alphas. The design of the surrogate model to predict Fitness Score ensures that our model can efficiently generate Alpha factors using gradient methods, even in scenarios characterized by a sparse search space. We also introduce a composite model capable of dynamically combining factor weights at each time slice, allowing the Mega-Alpha generated by the model to timely adjust its component factor and their weights to adapt to market fluctuations. Through extensive experiments, we have demonstrated that our framework can achieve better performance compared to previous methods of formulaic alpha models. In more realistic trading tests, our model consistently delivers higher profits and has already earning us excess returns during real money trading.